OpenStudy (anonymous):

(3x + 2y + z) + (6x - 5y - z) simplify

OpenStudy (anonymous):

(3x + 2y + z) + (6x - 5y -z) = (3x + 6x) + (2y - 5y) +(z -z)

OpenStudy (anonymous):

9x -3y

OpenStudy (anonymous):

(3x + 6x) + (2y - 5y) +(z -z) = 9x -3y + 0z

OpenStudy (akemi):

(3x + 2y + z) + (6x - 5y - z) = 3x+6x+2y-5y+z-z = 9x-3y+0z = 9x-3y